# disposition

### v2

@@ 2.2 Cones and their bases
$\operatorname{conv}\,K\text{ is well-based}\iff\operatorname{cl}(\operatorname{conv}\,K)\text{ is well-based}\iff\operatorname{int}\,K^{+}\neq\emptyset.$ (2.3)
Moreover, if $Y$ has finite dimension and $K$ is closed, then $K^{\#}=\operatorname{int}\,K^{+}$ (see [29, Th. 2.1(4), Rem. 2.6]), hence all the conditions involved in (2.2) and (2.3) are equivalent.
Example 2.4
Assume that $Y$ is a real normed space with the underlying norm $||\cdot||:Y\to\mathbb{R}$. For any $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$, a Bishop-Phelps cone (in the sense of Bishop and Phelps [2]) is defined by $C_{||\cdot||}(x^{*},\alpha):=\{y\in Y\mid x^{*}(y)\geq\alpha||y||\}$. It is known that Bishop-Phelps cones have a lot of useful properties and there are interesting applications in variational analysis and optimization (see, e.g., Ha and Jahn [36, 37], Phelps [55]). Any Bishop-Phelps cone $C_{||\cdot||}(x^{*},\alpha)$ is a closed, pointed, convex cone. If $||x^{*}||_{*}>\alpha$ (where $||\cdot||_{*}:Y^{*}\to\mathbb{R}$ denotes the dual norm of |||||$\cdot$||| | ⋅ | |), then $C_{||\cdot||}(x^{*},\alpha)$ is nontrivial and $C_{||\cdot||}^{>}(x^{*},\alpha):=\{y\in Y\mid x^{*}(y)>\alpha||y||\}=\operatorname{cor}\,C_{||\cdot||}(x^{*},\alpha)=\operatorname{int}\,C_{||\cdot||}(x^{*},\alpha)\neq\emptyset$. Note that for $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ with $||x^{*}||_{*}<\alpha$ we have $C_{||\cdot||}(x^{*},\alpha)=\{0\}$, while for $||x^{*}||_{*}\leq\alpha$ we have $C_{||\cdot||}^{>}(x^{*},\alpha)=\emptyset$. It is worth noting that every nontrivial, closed, convex cone $K\subseteq Y$ in a normed space $Y$ is representable as a Bishop-Phelps cone (i.e., there is a pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ and an equivalent norm $||\cdot||^{\prime}\sim||\cdot||$ such that $K=C_{||\cdot||^{\prime}}(x^{*},\alpha)$) if and only if $K$ is well-based (see the proof of Proposition 4.11). Note that any nontrivial, closed, pointed, convex cone $K$ in a finite dimensional normed space $Y$ is well-based, since in this setting $K$ is well-based if and only if $K^{\#}\neq\emptyset$ (taking into account Remark 2.3), and the latter condition $K^{\#}\neq\emptyset$ is true by the well-known Krein-Rutman theorem (see, e.g., [49, Th. 2.4.7]).
Moreover, we recall another (algebraic) base concept for cones in linear spaces (cf. Göpfert et al. [23, Def. 2.1.14], Jahn [42, Def. 1.10 (d)]).
Definition 2.5
@@
Remark 2.6
If $B$ is base in the sense of Definition 2.2 or Definition 2.5 for the nontrivial cone $K\subseteq Y$, then $K=\operatorname{\mathbb{R}}_{+}\cdot B$ and $K\setminus\{0\}=\operatorname{\mathbb{P}}\cdot B$.
@@
If $K$ is nontrivial and $\psi$ is positive on $K\setminus\{0\}$, then $B_{K}$ is called a normlike-base for $K$. If $\psi$ is a norm $||\cdot||:Y\to\operatorname{\mathbb{R}}$, then $B_{K}(||\cdot||)$ is called a norm-base for $K$. Note that $B_{K}(\psi)$ is a base in the sense of Definition 2.5, and if $\psi$ is continuous and $K$ is closed, then it is a base also in the sense of Definition 2.2.
2.3 Augmented dual cones and Bishop-Phelps type cones
@@
$\varphi_{x^{*},\alpha}(y):=x^{*}(y)+\alpha\psi(y)\quad\text{for all }y\in Y,$
as used in [28]. Note that Bishop-Phelps type (norm-linear) functions in a normed setting were studied for instance by García-Castaño, Melguizo-Padial and Parzanese [16], Gasimov (Kasimbeyli) [17], Günther, Khazayel and Tammer [29], Guo and Hang [31], Ha [35], Jahn [43], Kasimbeyli [45, 46], and Zaffaroni [59].
@@
Furthermore, it is easy to observe that $K^{a\#}\subseteq K^{a\circ}\cup K^{a+}$, and if $K$ is solid and $\psi$ is lower semicontinuous, then $K^{a\circ}\subseteq K^{a+}$.
In the following definition, we introduce two classes of functions that satisfy cone (interior) representation properties (cf. Ha [35, Sec. 3.1], Jahn [43, Th. 3.1]).
Definition 2.9
@@
Remark 2.10
It is known that several scalarizing functions in vector optimization (e.g., in the sense of Gerstewitz [20, 21]; Hiriart-Urruty [39] / Zaffaroni [58, Prop. 3.2]) admit cone representation properties. Moreover, for certain topological-linear spaces $Y$, which are preordered by a (solid) convex cone $K\subseteq Y$, one can state $-K$ representing (strictly $-K$ representing) functionals (see, e.g., Günther et al. [30, Sec. 4.1], Ha [35], and Jahn [43]).
@@
Remark that, in case $\psi=||\cdot||$ and $\alpha>0$, this reduces to the usual Bishop-Phelps cone (see Example 2.4). Moreover, the sets
@@
In the following lemmata, we collect some important properties of Bishop-Phelps type cones and functions. For the first lemma, the proof is straightforward and we omit it.
Lemma 2.11
@@
$3^{\circ}$ This assertion follows from Jahn [43, Prop. 3.6] taking into account $1^{\circ}$.
$4^{\circ}$ Of course, $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$ representing by $2^{\circ}$. Assume that $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$. By $1^{\circ}$ and $3^{\circ}$ we get $-S^{<}_{\psi}(x^{*},\alpha)=C^{>}_{\psi}(x^{*},\alpha)=$. Since $C_{\psi}(x^{*},\alpha)$ is a convex set with $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ it follows that $\operatorname{int}\,C_{\psi}(x^{*},\alpha)=$. Thus, we conclude $-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=-$ which shows that $\varphi_{x^{*},\alpha}$ is strictly $-C_{\psi}(x^{*},\alpha)$ representing.
Remark 2.14.
Note that the properties of $S^{\leq}_{\psi}(x^{*},\alpha)$ and $S^{<}_{\psi}(x^{*},\alpha)$, and the relationships between $S^{\leq}_{\psi}(x^{*},\alpha)$ and $S^{<}_{\psi}(x^{*},\alpha)$ follow immediately from the one related to $C_{\psi}(x^{*},\alpha)$ and $C^{>}_{\psi}(x^{*},\alpha)$. Furthermore, note that the conditions $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ (as given in the assumptions of Lemma 2.12 ($4^{\circ}$)) do not imply each other, as can be seen below:
Let us first give an example with $C^{>}_{\psi}(x^{*},\alpha)=\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$. According to [37, Ex. 2.1], for $Y:=\operatorname{\mathbb{R}}^{2}$, $\psi(\cdot):=||\cdot||_{1}$, $x^{*}:=(1,1)$ with $||x^{*}||_{\infty}=1,\alpha:=1$, one has $C_{\psi}(x^{*},1)=\{y\in\operatorname{\mathbb{R}}^{2}\mid x^{*}(y)=||y||_{1}\}=\operatorname{\mathbb{R}}^{2}_{+}$ and in this case $C^{>}_{\psi}(x^{*},\alpha)=\emptyset$, but $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$.
Next, we give an example with $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset$. Recall that, for a set $\Omega\subseteq Y$ in a real normed space $(Y,||\cdot||)$, one denotes the weak interior and weak closure of $\Omega$ by ${\mathrm{i}nt}_{w}\,\Omega$ and ${\mathrm{c}l}_{w}\,\Omega$, respectively. We have the next result:
Proposition 2.15.
@@
Proof 2.16.
On the contrary, assume that $\operatorname{int}_{w}\,C\neq\emptyset$. Take some $c\in\operatorname{int}_{w}\,C$, i.e., there is a weakly open neighbourhood $U(c)\subseteq Y$ around $c$ which satisfies $U(c)\subseteq C$. Since $Y$ has infinite dimension, the set $U(c)$ contains a line $c+\mathbb{R}\cdot v$ for some $v\in Y\setminus\{0\}$. This implies, for any $n\in\mathbb{N}$, we have $cn^{-1}\pm v=n^{-1}(c\pm nv)\in C$. Taking the limit $n\to\infty$, using the closedness of $C$, we conclude ±$\pm v\in C$, or equivalently, $v\in C\cap(-C)=\ell(C)=\{0\}$, a contradiction.
As a desired second example, we endow a real infinite dimensional (normed) space $Y$ with the weak topology, we put $\psi(\cdot):=||\cdot||$ and we assume $||x^{*}||^{*}>\alpha>0$. In this case, it is known that $C^{>}_{||\cdot||}(x^{*},\alpha)=\operatorname{int}_{||\cdot||}\,C_{||\cdot||}(x^{*},\alpha)\neq\emptyset$. Since $C_{||\cdot||}(x^{*},\alpha)$ is a closed, pointed cone in $(Y,||\cdot||)$ by Lemma 2.11 ($1^{\circ}$, $3^{\circ}$, $4^{\circ}$), we conclude that $\operatorname{int}_{w}\,C_{||\cdot||}(x^{*},\alpha)=\emptyset$ by Proposition 2.15.
Remark also that, if $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\psi$ is continuous, then $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ by Lemma 2.11 ($6^{\circ}$). In the second example from above, the weak upper semicontinuity of |||||$\cdot$||| | ⋅ | | is not satisfied, since $Y$ has infinite dimension (cf. [29, Rem. 2.2]).
2.4 Solution concepts in vector optimization
Given two real topological-linear spaces $X$ and $Y$, a seminorm $\psi:Y\to\mathbb{R}$, a nonempty feasible set $\Omega\subseteq X$, and a vector-valued objective function $f:X\rightarrow Y$, we consider the following vector optimization problem:
@@ 2.4 Solution concepts in vector optimization
Definition 2.17 (Efficiency).
A point $\overline{x}\in\Omega$ is said to be an efficient solution of ($P$) if there is no $x\in\Omega$ such that $f(x)\lneq_{K}f(\overline{x})$. The set of all Pareto efficient solutions of ($P$) is denoted by ${\mathrm{E}ff}(\Omega\mid f,K)$.
Definition 2.18 (Weak efficiency).
A point $\overline{x}\in\Omega$ is said to be a weakly efficient solution of ($P$) if there is no $x\in\Omega$ such that $f(x)<_{K}f(\overline{x})$. The set of all weakly Pareto efficient solutions of ($P$) is denoted by ${\mathrm{W}Eff}(\Omega\mid f,K)$.

### v3

@@ 2 Preliminaries
$\operatorname{conv}\,K\text{ is well-based}\iff\operatorname{cl}(\operatorname{conv}\,K)\text{ is well-based}\iff\operatorname{int}\,K^{+}\neq\emptyset.$ (4)
Moreover, if $Y$ has finite dimension and $K$ is closed, then $K^{\#}=\operatorname{int}\,K^{+}$ (see [15, Th. 2.1(4), Rem. 2.6]), hence all the conditions involved in (3) and (4) are equivalent.
Example 1 (Bishop-Phelps cone).
Assume that $Y$ is a real normed space with the underlying norm $||\cdot||:Y\to\mathbb{R}$. For any $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$, a Bishop-Phelps cone (in the sense of Bishop and Phelps [49]) is defined by
$C_{||\cdot||}(x^{*},\alpha):=\{y\in Y\mid x^{*}(y)\geq\alpha||y||\}.$
It is known that Bishop-Phelps cones have a lot of useful properties and there are interesting applications in variational analysis and optimization (see, e.g., Ha and Jahn [16, 17], Kasimbeyli [8], Phelps [50]). Any Bishop-Phelps cone $C_{||\cdot||}(x^{*},\alpha)$ is a closed, pointed, convex cone. If $||x^{*}||_{*}>\alpha$ (where $||\cdot||_{*}:Y^{*}\to\mathbb{R}$ denotes the dual norm of |||||$\cdot$||| | ⋅ | |), then $C_{||\cdot||}(x^{*},\alpha)$ is nontrivial and $C_{||\cdot||}^{>}(x^{*},\alpha):=\{y\in Y\mid x^{*}(y)>\alpha||y||\}=\operatorname{cor}\,C_{||\cdot||}(x^{*},\alpha)=\operatorname{int}\,C_{||\cdot||}(x^{*},\alpha)\neq\emptyset$. Note that for $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ with $||x^{*}||_{*}<\alpha$ we have $C_{||\cdot||}(x^{*},\alpha)=\{0\}$, while for $||x^{*}||_{*}\leq\alpha$ we have $C_{||\cdot||}^{>}(x^{*},\alpha)=\emptyset$.
It is worth noting that every nontrivial, closed, convex cone $K\subseteq Y$ in a normed space $Y$ is representable as a Bishop-Phelps cone (i.e., there is a pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ and an equivalent norm $||\cdot||^{\prime}\sim||\cdot||$ such that $K=C_{||\cdot||^{\prime}}(x^{*},\alpha)$) if and only if $K$ is well-based (see the proof of Proposition 23). Note that any nontrivial, closed, pointed, convex cone $K$ in a finite dimensional normed space $Y$ is well-based, since in this setting $K$ is well-based if and only if $K^{\#}\neq\emptyset$ (taking into account Remark 1), and the latter condition $K^{\#}\neq\emptyset$ is true by the well-known Krein-Rutman theorem (see, e.g., [42, Th. 2.4.7]). Further properties of Bishop-Phelps cones are studied in [51, Th. 4.5, Lem. 4.2], particularly another characterization for cones which are representable as a Bishop-Phelps cone (a so-called representation theorem). See also Lemma 20 and Remark 16.
Moreover, we recall another (algebraic) base concept for cones in linear spaces (cf. Göpfert et al. [21, Def. 2.1.14], Jahn [32, Def. 1.10 (d)] for the case of a convex cone).
Definition 2 (Algebraic base).
@@
Remark 2.
If $B$ is base in the sense of Definition 1 or Definition 2 for the nontrivial cone $K\subseteq Y$, then $K=\operatorname{\mathbb{R}}_{+}\cdot B$ and $K\setminus\{0\}=\operatorname{\mathbb{P}}\cdot B$.
@@
If $K$ is nontrivial and $\psi$ is positive on $K\setminus\{0\}$, then $B_{K}$ is called a normlike-base for $K$. If $\psi$ is a norm $||\cdot||:Y\to\operatorname{\mathbb{R}}$, then $B_{K}(||\cdot||)$ is called a norm-base for $K$. Note that $B_{K}(\psi)$ is a base in the sense of Definition 2, and if $\psi$ is continuous, then it is a base also in the sense of Definition 1.
Example 2 (Normlike-bases).
  • a)
    Consider a nontrivial cone $K\subseteq Y$ and some $x^{\ast}\in K^{\#}$. Define a seminorm $\psi$ by $\psi(x):=|x^{\ast}(x)|$ for all $x\in Y$. Notice that for dimension two or higher $\psi$ is not a norm. Since $x^{\ast}\in K^{\#}$, we have $\psi(x)>0$ for all $x\in K\setminus\{0\}$, hence $B_{K}(\psi)$ is a normlike-base for $K$. Notice that for $x^{\ast}\in K^{\#}$, we have
    $B_{K}(\psi)=K\cap\mathbb{S}_{\psi}=\{x\in K\mid|x^{\ast}(x)|=1\}=\{x\in K\mid x^{\ast}(x)=1\},$
    which is a usual type of bases for convex cones.
  • b)
    From [14, Lem. 2.16] we have the following results:
    Consider a function $\varphi:Y\to\mathbb{R}$, a function $\psi_{\max}:Y\to\mathbb{R}$ defined by
    $\psi_{\max}(y):=\max\{\varphi(y),\varphi(-y)\}\quad\text{ for all }y\in Y,$ (5)
    and assume that $K,A\subseteq Y$ are nontrivial cones. Then, the following assertions hold:
    • $1^{\circ}$
      If $\varphi$ is sublinear (i.e., $\varphi$ is positive homogeneous and subadditive), then $\psi_{\max}$ is a seminorm.
    • $2^{\circ}$
      Assume that $\varphi$ is $-K$ representing. Then, $\ell(K)=\{y\in Y\mid\psi_{\max}(y)\leq 0\}.$ Moreover, we have:
      $\psi_{\max}$ is nonnegative on $Y$ $\iff$ $\ell(K)=\{y\in Y\mid\psi_{\max}(y)=0\}$.
      $\psi_{\max}$ is positive on $Y\setminus\{0\}$ $\iff$ $K$ is pointed.
    • $3^{\circ}$
      Assume that $\varphi$ is sublinear and $-K$ representing. Then, $\psi_{\max}$ is a seminorm. Moreover, we have:
      $\ell(K)\cap A=\{0\}\iff$ $B_{A}(\psi_{\max})$ is a normlike-base for $A$.
    • $4^{\circ}$
      If $K$ is pointed, $\varphi$ is sublinear and $-K$ representing, then $\psi_{\max}$ is a norm, hence $B_{K}(\psi_{\max})$ and $B_{A}(\psi_{\max})$ are norm-bases for $K$ and $A$, respectively.
  • c)
    Assume that $A=\bigcup_{i=1}^{l}A_{i}$ for some nontrivial cones $A_{1},\ldots,A_{l}\subseteq Y$, and consider $y_{i}^{\ast}\in(A_{i})^{\#}$ for all $i=1,\ldots,l$. Define a seminorm $\psi$ (which is not necessarily a norm) by
    $\psi(x):=\max\{|x^{\ast}(x)|,|y_{1}^{\ast}(x)|,\ldots,|y_{l}^{\ast}(x)|\}\quad\text{for all }x\in Y,$
    or by
    $\psi(x):=|x^{\ast}(x)|+\sum_{i=1}^{l}\,|y_{i}^{\ast}(x)|\quad\text{for all }x\in Y.$
    Then, $\psi(x)>0$ for all $x\in(K\cup A)\setminus\{0\}$, hence $B_{K}(\psi)$ and $B_{A}(\psi)$ are normlike-bases for $K$ and $A$.
@@
$\varphi_{x^{*},\alpha}(y):=x^{*}(y)+\alpha\psi(y)\quad\text{for all }y\in Y,$ (6)
as used in [14]. Note that the seminorm $\psi$ must not have anything to do with the underlying topology of the real (locally convex) topological-linear space. Several examples of seminorms $\psi$, which are not necessarily norms, that can be used in the definition of the function $\varphi_{x^{*},\alpha}$ in (6) as well as in the subsequent study presented below, are discussed in [14, Sec. 2]. Bishop-Phelps type (norm-linear) functions in a normed setting (i.e., $\psi$ is given by the underlying norm of the space) were studied for instance by García-Castaño, Melguizo-Padial and Parzanese [1], Gasimov (Kasimbeyli) [6], Günther, Khazayel and Tammer [15], Guo and Hang [44], Ha [18], Jahn [20], Kasimbeyli [8, 9], and Zaffaroni [43].
@@
Furthermore, it is easy to observe that $K^{a\#}\subseteq K^{a\circ}\cup K^{a+}$, and if $K$ is solid and convex, and $\psi$ is lower semicontinuous, then $K^{a\circ}\subseteq K^{a+}$.
In the following definition, we introduce two classes of functions that satisfy cone (interior) representation properties (cf. Ha [18, Sec. 3.1], Jahn [20, Th. 3.1], see also Kasimbeyli [8, Th. 3.8 and 3.9]).
Definition 4 (Representing functionals).
@@
Remark that, in case $\psi=||\cdot||$ and $\alpha>0$, this reduces to the usual Bishop-Phelps cone (see Example 1). Moreover, the sets
@@
Example 3 (Bishop-Phelps type cones).
  • a)
    Assume that $Y$ has dimension two or higher, and consider the linear continuous functional $y^{*}\in Y^{*}\setminus\{0\}$ and the seminorm $\psi:=|y^{*}|$. Then, for the closed, convex cone $C:=\{y\in Y\mid y^{*}(y)\geq 0\}$ which is not pointed (and therefore not well-based if $Y$ is normed; see Example 1), we have
    $C=\{y\in Y\mid y^{*}(y)\geq|y^{*}(y)|\}=C_{\psi}(y^{*},1).$
    This means that the cone $C$ can be written as a Bishop-Phelps type cone (with $(y^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ and seminorm $\psi=|y^{*}|$). Note that $C$ cannot be written as a Bishop-Phelps cone (with $(y^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ and norm |||||$\cdot$||| | ⋅ | |), since any Bishop-Phelps cone is pointed, but $C$ is not.
  • b)
    Consider the function $\psi_{\max}:Y\to\mathbb{R}$ defined by (5) in Example 2(b), where $\varphi:Y\to\mathbb{R}$ is a sublinear function. For any $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$, the Bishop-Phelps type cone with respect to $\psi_{\max}$ is given by:
    $C_{\psi_{\max}}(x^{*},\alpha)$ =$=\{y\in Y\mid x^{*}(y)\geq\alpha\max\{\varphi(y),\varphi(-y)\}\}$
    =$=\{y\in Y\mid x^{*}(y)\geq\alpha\varphi(y)\text{ and }x^{*}(y)\geq\alpha\varphi(-y)\}\}.$
    Moreover, note that, for any function $\psi:Y\to\mathbb{R}$, the following statements are equivalent:
    • $1^{\circ}$
      $\psi$ is a seminorm.
    • $2^{\circ}$
      There is a sublinear function $\varphi:Y\to\mathbb{R}$ such that $\psi(y)=\max\{\varphi(y),\varphi(-y)\}$ for all $y\in Y$.
    Hence, the family of Bishop-Phelps type cones (for a fixed pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$) given by $\{C_{\psi_{\max}}(x^{*},\alpha)\mid\varphi:Y\to\mathbb{R}\text{ is sublinear}\}$ is equal to the family of Bishop-Phelps type cones given by $\{C_{\psi}(x^{*},\alpha)\mid\psi:Y\to\mathbb{R}\text{ is a seminorm}\}$.
  • c)
    Let $\mathcal{F}$ be the family of seminorms that generates the topology of the real locally convex space $Y$. Convex cones from the family of Bishop-Phelps type cones (for a fixed pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$) given by $\{C_{\psi}(x^{*},\alpha)\mid\psi\in\mathcal{F}\}$ could be of interest in order to define preorders in $Y$. In the case that the given space $Y$ is preordered but not partially ordered by a cone $K$, we get new ideas with considering Bishop-Phelps type type cones (since they are not necessarily pointed).
We collect next some important properties of Bishop-Phelps type cones and functions. For the first lemma, the proof is straightforward and we omit it.
Lemma 1.
@@
$3^{\circ}$ This assertion follows from Jahn [20, Prop. 3.6] taking into account $1^{\circ}$.
$4^{\circ}$ Of course, $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$ representing by $2^{\circ}$. Assume that $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$. By $1^{\circ}$ and $3^{\circ}$ we get $-S^{<}_{\psi}(x^{*},\alpha)=C^{>}_{\psi}(x^{*},\alpha)=\operatorname{cor}\,C_{\psi}(x^{*},\alpha)$. Since $C_{\psi}(x^{*},\alpha)$ is a convex set with $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ it follows that $\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\operatorname{cor}\,C_{\psi}(x^{*},\alpha)$. Thus, we conclude $-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=-\operatorname{cor}\,C_{\psi}(x^{*},\alpha)=-C^{>}_{\psi}(x^{*},\alpha)=S^{<}_{\psi}(x^{*},\alpha),$ which shows that $\varphi_{x^{*},\alpha}$ is strictly $-C_{\psi}(x^{*},\alpha)$ representing. ∎
Note that the properties of $S^{\leq}_{\psi}(x^{*},\alpha)$ and $S^{<}_{\psi}(x^{*},\alpha)$, and the relationships between $S^{\leq}_{\psi}(x^{*},\alpha)$ and $S^{<}_{\psi}(x^{*},\alpha)$ follow immediately from the one related to $C_{\psi}(x^{*},\alpha)$ and $C^{>}_{\psi}(x^{*},\alpha)$. Furthermore, note that the conditions $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ (as given in the assumptions of Lemma 2 ($4^{\circ}$)) do not imply each other, as can be seen below:
Let us first give an example with $C^{>}_{\psi}(x^{*},\alpha)=\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$. According to [17, Ex. 2.1], for $Y:=\operatorname{\mathbb{R}}^{2}$, $\psi(\cdot):=||\cdot||_{1}$, $x^{*}:=(1,1)$ with $||x^{*}||_{\infty}=1,\alpha:=1$, one has $C_{\psi}(x^{*},1)=\{y\in\operatorname{\mathbb{R}}^{2}\mid x^{*}(y)=||y||_{1}\}=\operatorname{\mathbb{R}}^{2}_{+}$ and in this case $C^{>}_{\psi}(x^{*},\alpha)=\emptyset$, but $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$.
Next, we give an example with $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset$. We endow a real infinite dimensional (normed) space $Y$ with the weak topology, we put $\psi(\cdot):=||\cdot||$ and we assume $||x^{*}||^{*}>\alpha>0$. In this case, it is known that $C^{>}_{||\cdot||}(x^{*},\alpha)=\operatorname{int}_{||\cdot||}\,C_{||\cdot||}(x^{*},\alpha)\neq\emptyset$. Since $C_{||\cdot||}(x^{*},\alpha)$ is a closed, pointed cone in $(Y,||\cdot||)$ by Lemma 1 ($1^{\circ}$, $3^{\circ}$, $4^{\circ}$), we conclude that $\operatorname{int}_{w}\,C_{||\cdot||}(x^{*},\alpha)=\emptyset$, as the following proposition shows. Recall that, for a set $\Omega\subseteq Y$ in a real normed space $(Y,||\cdot||)$, one denotes the weak interior and weak closure of $\Omega$ by ${\rm int}_{w}\,\Omega$ and ${\rm cl}_{w}\,\Omega$, respectively.
Proposition 3.
@@
Proof.
On the contrary, assume that $\operatorname{int}_{w}\,C\neq\emptyset$. Take some $c\in\operatorname{int}_{w}\,C$, i.e., there is a weakly open neighbourhood $U(c)\subseteq Y$ around $c$ which satisfies $U(c)\subseteq C$. Since $Y$ has infinite dimension, the set $U(c)$ contains a line $c+\mathbb{R}\cdot v$ for some $v\in Y\setminus\{0\}$. This implies, for any $n\in\mathbb{N}$, we have $cn^{-1}\pm v=n^{-1}(c\pm nv)\in C$. Taking the limit $n\to\infty$, using the closedness of $C$, we conclude ±$\pm v\in C$, or equivalently, $v\in C\cap(-C)=\ell(C)=\{0\}$, a contradiction. ∎
Remark also that, if $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\psi$ is continuous, then $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ by Lemma 1 ($6^{\circ}$). In the second example from above, the weak upper semicontinuity of |||||$\cdot$||| | ⋅ | | is not satisfied, since $Y$ has infinite dimension (cf. [15, Rem. 2.2]).
In what follows, we will briefly introduce different solution concepts in vector optimization. Given two real topological-linear spaces $X$ and $Y$, a nonempty feasible set $\Omega\subseteq X$, and a vector-valued objective function $f:X\rightarrow Y$, we consider the following vector optimization problem:
@@
Definition 5 (Efficiency).
A point $\overline{x}\in\Omega$ is said to be an efficient solution of ($P$) if there is no $x\in\Omega$ such that $f(x)\lneq_{K}f(\overline{x})$. The set of all efficient solutions of ($P$) is denoted by ${\rm Eff}(\Omega\mid f,K)$.
Definition 6 (Weak efficiency).
A point $\overline{x}\in\Omega$ is said to be a weakly efficient solution of ($P$) if there is no $x\in\Omega$ such that $f(x)<_{K}f(\overline{x})$. The set of all weakly efficient solutions of ($P$) is denoted by ${\rm WEff}(\Omega\mid f,K)$.